Use Lagrange multipliers to find the point on the plane x − 2y + 3z = 6 that is closest to the point (0, 3, 5)
Answer #1

Let (2, 4,7) be required point let f- ( distance from cal, y, z) ló (0,3,5)) f = x2 +ly-3)2+(7-5) and g = 3-2y+37=6 . Stepis20) = -- (5-2) I 3 ju - - 39 + 6 67 (બ, ૨, ૩ ( (ભ , 9, ) = ક )